OpenStudy (anonymous):

If cos Θ = -2/5 and tan Θ > 0, what is the value of sin Θ? @Hero

zepdrix (zepdrix):

\(\Large\rm \cos\theta\) is negative. \(\Large\rm \tan\theta\) is positive. So what quadrant does that put our angle \(\Large\rm\theta\) in?

OpenStudy (anonymous):

Quadrant II?

OpenStudy (anonymous):

my bad, quadrant IV

zepdrix (zepdrix):

There's a uhhh weird acronym to help remember.. Lemme see if I can recall.. It's like.. "All Students Take Calculus" or maybe you've learned something similar :P I dunno. ASTC, All are positive in Q1, Sine positive in Q2, Tangent positive in Q3, Cosine positive in Q4. So if have tangent positive, but not cosine, we must be in Q3

zepdrix (zepdrix):

So draw a triangle in Q3,

zepdrix (zepdrix):